# Tomographic reconstruction of a disease transmission landscape via GPS recorded random paths

### v3

@@ 5 Conclusion
We introduce an innovative model that extends tomographic methodologies to estimate landscape susceptibility to various risks using GPS data, in a manner analogous to Positron Emission Tomography (PET). By adapting tomographic principles to interpret spatial data from GPS signals, our model indirectly assesses landscape-related risks, broadening tomography’s application beyond medical imaging. Our approach uniquely predicts disease transmission risk without requiring detailed knowledge of vector ecology or environmental conditions, distinguishing it from traditional models that rely on extensive host movement and vector presence data. The sample size required and the time series length will depend on the geographical extent, ecology of the disease and movement ecology of the animals.
In addition to developing this novel approach, we propose a solution based on Total Variation with a selection of its smoothing parameter. We validate our approach with WTD data, observing alignment with known EHDV risks and vector ecology, while providing additional insights to investigate. Comparative analyses reveal our model’s performance exceeds that of conventional spatial statistical and inverse problem-solving approaches, underscoring its robustness and adaptability. While individual recapture data is required—posing challenges for certain species—the model is applicable in contexts where movement tracking is feasible, such as avian malaria research.
@@
Appendix A Proof of Property 2.1

### v4

@@ 7 Conclusion
We introduce an innovative model that extends tomographic methodologies to estimate landscape susceptibility to various risks using GPS data, in a manner analogous to Positron Emission Tomography (PET). By adapting tomographic principles to interpret spatial data from GPS signals, our model indirectly assesses landscape-related risks, broadening tomography’s application beyond medical imaging. Our approach uniquely predicts disease transmission risk without requiring detailed knowledge of vector ecology or environmental conditions, distinguishing it from traditional models that rely on extensive host movement and vector presence data. The sample size required and the time series length depend on the geographical extent, ecology of the disease and movement ecology of the animals.
We validate our approach with WTD data, observing alignment with known EHDV risks and vector ecology, while providing additional insights to investigate. Comparative analyses reveal our model’s performance exceeds that of conventional spatial statistical and inverse problem-solving approaches, underscoring its robustness and adaptability. While individual recapture data is required—posing challenges for certain species—the model is applicable in contexts where movement tracking is feasible, such as avian malaria research.
In addition to developing this novel approach, we propose a selection of the total variation regularization parameter suitable for high-dimensional linear inverse problems.
{acks}
[Acknowledgments] The authors would like to thank the University of St. Gallen and the Leading House for the Latin American Region for their support through the Swiss-Latin America Seed Money Grant. First author acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC); through grant DGECR-2022-04531. Data and funding for the field studies was supported by University of Florida, Institute for Food and Agricultural Sciences, Cervidae Health Research Initiative, funded through the Florida State Legislature, grant number LBR2199. The owner of the private deer ranch at the time of deer tracking enabled this study and we thank the ranch managers for their extensive logistical support. We also thank colleagues at the Florida Medical Entomology Laboratory for collecting the entomological data.
@@
Appendix A Proof of Property 3.1